OpenStudy (lgbasallote):

A substance, \(\text A_2\text B\), has the composition by mass of 60% A and 40% B. What is the composition of \(\text{AB}_2\) by mass?

OpenStudy (australopithecus):

yeah its wrong I need to rethink it :S

OpenStudy (australopithecus):

This is definitely a system of equations problem though

OpenStudy (australopithecus):

because A and B are constant, we just need to think of two equations we can use to solve the problem

OpenStudy (lgbasallote):

isnt it a mass percent problem?

OpenStudy (australopithecus):

Yeah it is

OpenStudy (australopithecus):

We Know \[\frac{Molecular Mass_{2A}}{Molecular Mass_{A_2B}} = 0.6\] \[\frac{Molecular Mass_{B}}{Molecular Mass_{A_2B}} = 0.4\]

OpenStudy (australopithecus):

Just use system of equations to solve the problem, isolate the molecular mass of A2B and sub it into the other equation

OpenStudy (australopithecus):

then you can solve for B and 2A

OpenStudy (australopithecus):

then just divide 2A by 2 to find the molecular mass of A

OpenStudy (lgbasallote):

lol you confuse me

OpenStudy (lgbasallote):

you bombard me with words and i just cant follow

OpenStudy (australopithecus):

Yeah sorry I'm always all over the place

OpenStudy (australopithecus):

do you not understand my method?

OpenStudy (lgbasallote):

*understand* yes

OpenStudy (australopithecus):

Do you not agree with it?

OpenStudy (australopithecus):

wait never mind I need to think about this some more my method doesn't make sense

OpenStudy (lgbasallote):

i have absolutely have no idea to even have a position whether to agree or disagree

OpenStudy (australopithecus):

lol

OpenStudy (lgbasallote):

also i didnt get what you said so i cant really agree or disagree

OpenStudy (australopithecus):

fair enough

OpenStudy (australopithecus):

Sorry I don't think I have come across a problem such as this, also it has been a long time since i have had to do such problem

OpenStudy (lgbasallote):

aww :(

OpenStudy (australopithecus):

Wait I think I might have it :L \[\frac{MolecularMass_{2A}}{0.6} = MolecularMass_{A_{2}B}\] \[\frac{MolecularMass_{B}}{0.4} = MolecularMass_{A_2B}\] \[\frac{MolecularMass_{B}}{0.4} = \frac{MolecularMass_{2A}}{0.6}\] = \[(0.6)MolecularMass_{B} = (0.4)MolecularMass_{2A}\] = \[\frac{(0.6)MolecularMass_{B}}{(0.4)} = MolecularMass_{2A}\] Now you have Them molecular mass of B in terms of 2A :)

OpenStudy (australopithecus):

the* not them

OpenStudy (australopithecus):

Blah I might be way off on my solution give me a second if I'm I recommend you repost the question and I appologize for taking up your time

OpenStudy (australopithecus):

Yeah I dont know how to solve this best to just reask the question I bet the solution is easy :S

OpenStudy (anonymous):

whats the ans??

OpenStudy (anonymous):

Out of 100g two A atoms weigh 60g(1 A atom weighs 30g) and 1 B atom weighs 40g. In the sec molecule, there is 1 A atom of 30g and 2 B atoms of 80g. So the mass percent of A is 30*100/110=27 and B that of B is 73(approx.)

OpenStudy (unklerhaukus):

\[A_2B=100\%= 60\% A + 40\% B\] \[A_2=60\%A_2B\qquad\qquad B=40\%A_2B\]\[A=30\%A_2B\]__________ \[AB_2=30\%A_2B+2\times40\%A_2B=110\%A_2B\] \[A=30\%A_2B\times \frac{100\%AB_2}{110\%A_2B}=27.\dot2\dot7\%AB_2\] \[B=2\times 40\%A_2B\times\frac{100\%AB_2}{110\%A_2B}=72.\dot7\dot2\%AB_2\]